OpenStudy (anonymous):

f(x) = (3^x+2 ) -5 find f^-1 (x)

OpenStudy (anonymous):

go to www.wolframalpha.com. they will show u steps along with answer :)

OpenStudy (zzr0ck3r):

\(y=3^x+2\implies y-2=3^x\implies \log_3(y-2)=x\) so the inverse function is \(f(x) = \log_3(x-2 )\)

OpenStudy (anonymous):

thanks. the wolfram thing didn't make any sense =.=